Michigan Co. is expected to pay a dividend of $1.50 per share one year from now. The dividends are expected to grow at 15% per year for the next six years (years 1-6) and then grow 5% per year thereafter (from year 7 and beyond). The discount rate is 18%. What is the current value of the company’s stock? Please round your answer to the nearest penny.
Answer #1

D1 = $1.5

D2= $1.7250 (1.5*1.15)

D3 = 1.9838 (1.725*1.15)

D4=$2.2813 (1.9838*1.15)

D5 =$2.6235 ($2.2813*1.15)

D6 =$3.017 (2.6235*1.15)

D7 = 3.1675 (3.017*1.05)

P6 = D7/ Re - g

= 3.1675/0.18 - 0.05

=$24.3681

So, the current value of the company's stock is :

1.5/1.18 + 1.725/1.18^2 + 1.9838/1.18^3 + 2.2813/1.18^4 + 2.6235/1.18^5 + 3.017/1.18^6

=1.2712 + 1.2389 + 1.2074 + 1.1767 +1.1468 + 1.1176

=$5.9819

current value of stock is = $5.9819 + $24.3681/1.18^6

=5.9819 + 9.0267 = $15.0086

therefore. the current value of Michigan stock is $15.